# Gx2Mol: De Novo Generation of Hit-like Molecules from Gene Expression Profiles via Deep Learning

### v1

@@ 4.5 Case Studies
Figure 12: Assessing therapeutic impact: A comparison of Tanimoto coefficients and molecular structures between Gx2Mol and DRAGONET baseline model for newly generated molecules.
Figure 11 illustrates the data processing of a gene expression profile for the generation of molecules with therapeutic effects on a disease. First, as shown in Figure 11 (A), a disease-specific gene expression profile is constructed by averaging the gene expression profiles of patients with a certain disease. Then, a gene expression profile that is inversely correlated with the disease-specific gene expression profile is constructed and defined as the disease reversal profile, as shown in Figure 11 (B). Finally, the disease reversal profile is used as an input to the proposed Gx2Mol model to generate molecules with therapeutic effects (Figure 11 (C)). The disease-specific gene expression profiles were obtained from the CREEDS database for patients with three diseases, i.e., gastric cancer, atopic dermatitis, and Alzheimer’s disease.
We examined the validity of the newly generated molecules by comparing the newly generated molecular structures with those of the approved drugs. If the newly generated molecules are meaningful, the newly generated molecules should be structurally similar to the approved drugs of each disease to some extent. We calculated the structural similarity using Tanimoto coefficients. Figure 12 illustrates the Tanimoto coefficients between approved drugs and newly generated molecules, comparing the results obtained from the SOTA DRAGONET [64] and our proposed Gx2Mol model, for each of the three diseases. Our proposed Gx2Mol model surpassed the SOTA DRAGONET in hit-like molecule generation for three diseases. Gx2Mol exhibited improved Tanimoto coefficients to approved drugs, reaching 0.58, 0.60, and 0.53 for gastric cancer, Alzheimer’s disease, and atopic dermatitis. These improvements amounted to 13.7%, 130.8%, and 82.8% for these three diseases. Additionally, fluorouracil (D04197) can be used in the treatment of liver metastases from gastrointestinal adenocarcinomas and also in the palliative treatment of liver and gastrointestinal cancers [65]. When using the disease reversal profile of gastric cancer patients, the Tanimoto coefficient between the molecule generated by the proposed Gx2Mol model and fluorouracil was the largest. The Tanimoto coefficient of the Gx2Mol model generated molecule with floxuridine was maximum using the disease reversal profile of gastric cancer patients. These results suggest that the generated molecules effectively capture the structural features of approved anti-gastric cancer drugs. In addition, the molecules generated for the other two diseases demonstrate structural features that are similar to those of the approved drugs. As a result, the molecules generated using the proposed Gx2Mol model have potential drug-like properties.
@@ 5 Conclusion
This study introduced the Gx2Mol model, designing to generate potential chemical structures of hit-like molecules from gene expression profiles using deep learning techniques. In the training phase, the Gx2Mol model first employed a VAE for feature extraction from high-dimensional gene expression profiles, and then the low-dimensional extracted features guided the generation of syntactically valid SMILES strings. In the generation phase, the VAE encoder served as the sole feature extractor, seamlessly combined with the generator to facilitate the generation of hit-like molecules. The results demonstrated the effectiveness of Gx2Mol in generating hit-like molecules from gene expression profiles. Additionally, a case study illustrates the model’s ability to generate potential chemical structures for therapeutic drugs related to gastric cancer, stress dermatitis, and Alzheimer’s disease using patients’ disease reversal profiles.
This study has a primary limitation. Since LSTMs are frequently employed in auto-regressive generation tasks, wherein the token at the next time step is generated based on the token at the current time step, there exists a potential constraint on the diversity of generated molecules when using LSTMs as generators. In future research, we aim to explore strategies to enhance the diversity of molecular generation within the proposed Gx2Mol model. Furthermore, the envisaged application of the Gx2Mol model involves integration into practical AI systems to assist chemists in generating diverse drug candidate hit-like molecules tailored for various diseases. This integration is anticipated to leverage the strengths of the Gx2Mol model and contribute to the advancement of drug discovery processes.

### v2

@@ 4.5 Case Studies
Figure 12: Assessing therapeutic impact: A comparison of Tanimoto coefficients and molecular structures between HVL2Mol and DRAGONET baseline model for newly generated molecules.
Figure 11 illustrates the data processing of a gene expression profile for the generation of molecules with therapeutic effects on a disease. First, as shown in Figure 11 (A), a disease-specific gene expression profile is constructed by averaging the gene expression profiles of patients with a certain disease. Then, a gene expression profile that is inversely correlated with the disease-specific gene expression profile is constructed and defined as the disease reversal profile, as shown in Figure 11 (B). Finally, the disease reversal profile is used as an input to the proposed HVL2Mol model to generate molecules with therapeutic effects (Figure 11 (C)). The disease-specific gene expression profiles were obtained from the CREEDS database for patients with three diseases, i.e., gastric cancer, atopic dermatitis, and Alzheimer’s disease.
We examined the validity of the newly generated molecules by comparing the newly generated molecular structures with those of the approved drugs. If the newly generated molecules are meaningful, the newly generated molecules should be structurally similar to the approved drugs of each disease to some extent. We calculated the structural similarity using Tanimoto coefficients. Figure 12 illustrates the Tanimoto coefficients between approved drugs and newly generated molecules, comparing the results obtained from the SOTA DRAGONET [64] and our proposed HVL2Mol model, for each of the three diseases. Our proposed HVL2Mol model surpassed the SOTA DRAGONET in hit-like molecule generation for three diseases. HVL2Mol exhibited improved Tanimoto coefficients to approved drugs, reaching 0.58, 0.60, and 0.53 for gastric cancer, Alzheimer’s disease, and atopic dermatitis. These improvements amounted to 13.7%, 130.8%, and 82.8% for these three diseases. Additionally, fluorouracil (D04197) can be used in the treatment of liver metastases from gastrointestinal adenocarcinomas and also in the palliative treatment of liver and gastrointestinal cancers [65]. When using the disease reversal profile of gastric cancer patients, the Tanimoto coefficient between the molecule generated by the proposed HVL2Mol model and fluorouracil was the largest. The Tanimoto coefficient of the HVL2Mol model generated molecule with floxuridine was maximum using the disease reversal profile of gastric cancer patients. These results suggest that the generated molecules effectively capture the structural features of approved anti-gastric cancer drugs. In addition, the molecules generated for the other two diseases demonstrate structural features that are similar to those of the approved drugs. As a result, the molecules generated using the proposed HVL2Mol model have potential drug-like properties.
@@ 5 Conclusion
This study introduced the HVL2Mol model, designing to generate potential chemical structures of hit-like molecules from gene expression profiles using deep learning techniques. In the training phase, the HVL2Mol model first employed a VAE for feature extraction from high-dimensional gene expression profiles, and then the low-dimensional extracted features guided the generation of syntactically valid SMILES strings. In the generation phase, the VAE encoder served as the sole feature extractor, seamlessly combined with the generator to facilitate the generation of hit-like molecules. The results demonstrated the effectiveness of HVL2Mol in generating hit-like molecules from gene expression profiles. Additionally, a case study illustrates the model’s ability to generate potential chemical structures for therapeutic drugs related to gastric cancer, stress dermatitis, and Alzheimer’s disease using patients’ disease reversal profiles.
This study has a primary limitation. Since LSTMs are frequently employed in auto-regressive generation tasks, wherein the token at the next time step is generated based on the token at the current time step, there exists a potential constraint on the diversity of generated molecules when using LSTMs as generators. In future research, we aim to explore strategies to enhance the diversity of molecular generation within the proposed HVL2Mol model. Furthermore, the envisaged application of the HVL2Mol model involves integration into practical AI systems to assist chemists in generating diverse drug candidate hit-like molecules tailored for various diseases. This integration is anticipated to leverage the strengths of the HVL2Mol model and contribute to the advancement of drug discovery processes.